OpenStudy (precal):

possible vertical tangent problem f(x)=(ln(cosx))/x^2 find the derivative at x=pi

OpenStudy (precal):

ok I worked it out, I need someone to doublecheck my conclusion

OpenStudy (precal):

f'(pi) is undefined and therefore f(x) has a vertical tangent?

OpenStudy (sidsiddhartha):

at x=pi f'(pi)=-0.202642i

OpenStudy (amoodarya):

\[\cos x>0 \rightarrow Df(x): -\frac{ \pi }{2 }+2k \pi < x<+\frac{ \pi }{2 }+2k \pi -\left\{ 0 \right\}\]

OpenStudy (amoodarya):